# Reframing the Relationship in Out-of-Distribution Detection

### v1

@@ 4.3 Ablation study
We conducted a comparative analysis of the performance of CMA across various models [59], with the results presented in Table  5. The findings indicate that as the backbone model’s performance improves, the effectiveness of CMA also intensifies, resulting in superior OOD detection capabilities. Furthermore, this study underscores the fact that our approach is not reliant on a single model, but can be effectively integrated across diverse models, demonstrating remarkable scalability.
5 Discussion
Figure 4: Impact curve of $k$ on the performance of CMA. $k=\frac{\text{number of Agents}}{\text{number of ID Labels}}$
Figure 5: Comparison with different Agents.
The number of Agents.
In section 3, we present the score calculation formula $\mathcal{S}_{\text{CMA}}(\mathbf{x}^{\prime};\mathcal{Y}_{in},\mathcal{Y}_{ntc},\mathcal{T},\mathcal{I})$ for CMA, where we set the number of Agents to $N$, aligning with the number of ID categories. This number was determined based on extensive experiments, with $k$ set to $k=\frac{M}{N}$, where $M$ is the number of Agents.Concurrently, the formula for calculating the score has been modified to $\mathcal{S}_{\text{CMA}}^{\prime}(\mathbf{x}^{\prime};\mathcal{Y}_{in},\mathcal{Y}_{ntc},\mathcal{T},\mathcal{I})=\frac{\exp(\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{\hat{y}})/\tau)}{\sum_{i=1}^{N}\exp(\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{i})/\tau)+\sum_{i=1}^{M}\exp(\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{i})/\tau)}$, where $\hat{y}=\mathop{\arg\max}\limits_{1\leq i\leq N}\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{i})$. Experiments were conducted for various values of $k$. Figure 4 illustrates the results of our experiments conducted on ImageNet-1k. The optimal value of $k$ was identified as 1. Additionally, it was observed that an excessive number of Agents does not enhance the overall performance of CMA, whereas a insufficient number of Agents can adversely affect it. This finding illuminates the rationality of the vector triangle relationship. An insufficient number of Agents fails to exert the same impact on OOD data as it does on ID categories, leading to structural disarray and performance degradation. Once the number of Agents reaches a certain threshold, a stable structure is established, after which increasing the number of Agents does not significantly improve performance.
Performance of different Agents.
Figure 5 illustrates the impact of different neutral terms on the OOD detection performance of agents. It is evident that the impact of different agents on performance is substantial. On average, the best and worst outcomes in terms of FPR95 and AUROC can differ by 19.95% and 4.50%, respectively. From another perspective, certain agents excel in specific contexts, such as "a large photo of the scene we live in." While their overall performance across multiple environments lags significantly behind the optimal level, they achieve their best results when using the SUN dataset [77] as OOD data. On one hand, this highlights a potential limitation of CMA, where different agents can yield disparate outcomes. On the other hand, it underscores the significant potential of CMA, as it can tailor specific agents to suit diverse environments, thereby enhancing performance.
6 Related Works
@@
Visual Out-of-distribution (OOD) detection [79, 44, 80], a significant area in machine learning systems and computer vision, involves discerning images devoid of specific semantic information. Following a prolonged period of development, this field has evolved into two broad categories: single-modality OOD detection methods and multimodal OOD detection approaches that leverage semantic information. For the former, D Hendrycks et al. [24] outlines the fundamental paradigm within neural networks, which has been the foundation for numerous subsequent developments, including output-based methods [4, 15, 17, 23, 26, 28, 45, 64], density-based methods [1, 88, 12, 63, 56], distance-based methods [38, 49, 62, 65], reconstruction-based methods [39, 14, 87, 82], and gradient-based methods [42, 29, 31]. Some works [53, 20, 5] have also provided enhanced theoretical analyses for OOD detection. Recent studies have sought to reexamine and address the OOD detection problem from various perspectives. Ammar et al. [3] proposes utilizing "neural collapse" and the geometric properties of principal component spaces to identify OOD data. Lu et al. [46] models each class with multiple prototypes to learn more compact sample embeddings, thereby enhancing OOD detection capabilities.
@@
As multimodal models emerge as a new trend, multimodal pretrained models [59, 32, 78] are increasingly being introduced into other domains [81, 40, 35, 60, 43, 19]. The idea of utilizing large-scale pretrained multimodal models that rely solely on outlier class names without any accompanying images was first proposed in Fort et al. [21]. ZOC [18] employs an extended model to generate candidate unknown class names for each test sample. MCM[48], for the first time, abandons the traditional reliance on sample assumptions, starting from a language-vision representation and using a temperature scaling strategy and the maximum predicted softmax value as the out-of-distribution (OOD) score, establishing a paradigm for CLIP-based OOD detection. Building on MCM, GL-MCM [51] enhances the model by aligning the global and local visual-textual features of CLIP [59]. LoCoOp[52] proposes a few-shot OOD detection method within the framework of prompt learning. LSN [54], NegPrompt [41], and NegLabel [33] all enhance the effectiveness of OOD detection by introducing negative prompts and utilizing methods such as prompt learning and sampling. Our approach, CMA, is rooted in MCM and offers a fresh perspective on CLIP-based OOD detection from the language-vision representation standpoint. It is a post-hoc detection method that obviates the need for learning and external data.
7 Conclusion
Appendix A Experimental Details
A.1 Software and Hardware
@@
A.2 Datasets
ImageNet-10 We use the ImageNet-10 in [48] that mimics the class distribution of CIFAR-10 but with high-resolution images. It contains the following categories (with class ID): warplane (n04552348), sports car (n04285008), brambling bird, (n01530575), Siamese cat (n02123597), antelope (n02422699), Swiss mountain dog (n02107574), bull frog (n01641577), garbage truck (n03417042), horse (n02389026), container ship (n03095699).
ImageNet-20 We use the ImageNet-20 in [48], which consists of 20 classes semantically similar to ImageNet-10 (e.g. dog (ID) vs. wolf (OOD)). It contains the following categories: sailboat (n04147183), canoe (n02951358), balloon (n02782093), tank (n04389033), missile (n03773504), bullet train (n02917067), starfish (n02317335), spotted salamander (n01632458), common newt (n01630670), zebra (n01631663), frilled lizard (n02391049), green lizard (n01693334), African crocodile (n01697457), Arctic fox (n02120079), timber wolf (n02114367), brown bear (n02132136), moped (n03785016), steam locomotive (n04310018), space shuttle (n04266014), snowmobile (n04252077).
ImageNet-100 We use the ImageNet-100 in [48], which is curated from ImageNet-1k.
@@ A.2 Datasets
A.3 Sources of model checkpoints
@@ A.4 Hyperparameters
Appendix B ID Classification Accuracy

### v2

@@ D.2 Comparative performance across models
We conducted a comparative analysis of the performance of CMA across various models (Radford et al. 2021), with the results presented in Table  5. The findings indicate that as the backbone model’s performance improves, the effectiveness of CMA also intensifies, resulting in superior OOD detection capabilities. Furthermore, this study underscores the fact that our approach is not reliant on a single model, but can be effectively integrated across diverse models, demonstrating remarkable scalability.
Appendix E Related Works
@@
Visual Out-of-distribution (OOD) detection (Yang et al. 2021; Liu et al. 2021; Yang et al. 2022), a significant area in machine learning systems and computer vision, involves discerning images devoid of specific semantic information. Following a prolonged period of development, this field has evolved into two broad categories: single-modality OOD detection methods and multimodal OOD detection approaches that leverage semantic information. For the former, D Hendrycks et al. (Hendrycks and Gimpel 2016) outlines the fundamental paradigm within neural networks, which has been the foundation for numerous subsequent developments, including output-based methods (Bendale and Boult 2016; DeVries and Taylor 2018; Du et al. 2022; Hein, Andriushchenko, and Bitterwolf 2019; Hsu et al. 2020; Huang and Li 2021; Liu et al. 2020; Sun and Li 2022), density-based methods (Abati et al. 2019; Zong et al. 2018; Deecke et al. 2019; Sabokrou et al. 2018; Pidhorskyi, Almohsen, and Doretto 2018), distance-based methods (Lee et al. 2018; Ming et al. 2022b; Ren et al. 2021; Sun et al. 2022), reconstruction-based methods (Li et al. 2023a; Denouden et al. 2018; Zhou 2022; Yang, Gao, and Xu 2022), and gradient-based methods (Liang, Li, and Srikant 2017; Huang, Geng, and Li 2021; Igoe et al. 2022). Some works (Morteza and Li 2022; Fang et al. 2022; Bitterwolf et al. 2022) have also provided enhanced theoretical analyses for OOD detection. Recent studies have sought to reexamine and address the OOD detection problem from various perspectives. Ammar et al. (Ammar et al. 2024) proposes utilizing ”neural collapse” and the geometric properties of principal component spaces to identify OOD data. Lu et al. (Lu et al. 2023) models each class with multiple prototypes to learn more compact sample embeddings, thereby enhancing OOD detection capabilities.
@@
As multimodal models emerge as a new trend, multimodal pretrained models (Radford et al. 2021; Jia et al. 2021; Xu et al. 2022) are increasingly being introduced into other domains (Yang et al. 2023; Li et al. 2023b; Kirillov et al. 2023; Radford et al. 2023; Lin et al. 2023; Fang et al. 2023). The idea of utilizing large-scale pretrained multimodal models that rely solely on outlier class names without any accompanying images was first proposed in Fort et al. (Fort, Ren, and Lakshminarayanan 2021). ZOC (Esmaeilpour et al. 2022) employs an extended model to generate candidate unknown class names for each test sample. MCM(Ming et al. 2022a), for the first time, abandons the traditional reliance on sample assumptions, starting from a language-vision representation and using a temperature scaling strategy and the maximum predicted softmax value as the out-of-distribution (OOD) score, establishing a paradigm for CLIP-based OOD detection. Building on MCM, GL-MCM (Miyai et al. 2023) enhances the model by aligning the global and local visual-textual features of CLIP (Radford et al. 2021). LoCoOp(Miyai et al. 2024) proposes a few-shot OOD detection method within the framework of prompt learning. LSN (Nie et al. 2023), NegPrompt (Li et al. 2024), and NegLabel (Jiang et al. 2024) all enhance the effectiveness of OOD detection by introducing negative prompts and utilizing methods such as prompt learning and sampling. Our approach, CMA, is rooted in MCM and offers a fresh perspective on CLIP-based OOD detection from the language-vision representation standpoint. It is a post-hoc detection method that obviates the need for learning and external data.
Appendix F ID Classification Accuracy
Appendix G Limitation
Our proposed method, CMA, reconstructs the relationships in OOD detection by incorporating neutral prompts as agents, thereby enhancing the capability of OOD detection. However, the integration of external information is a double-edged sword. On one hand, if the neutral prompts share similar semantics with the ID labels, it can compromise the model’s ID detection capabilities. On the other hand, external information struggles to mitigate the impact of OOD images that are too similar to the ID labels, as it does not directly augment the model’s intrinsic abilities. Additionally, the challenge remains in obtaining neutral prompts that align more closely with the requirements of the ID labels without the need for training, necessitating further in-depth research.
Appendix H Experimental Details
H.1 Software and Hardware
@@
H.2 Datasets
ImageNet-10 We use the ImageNet-10 in (Ming et al. 2022a) that mimics the class distribution of CIFAR-10 but with high-resolution images. It contains the following categories (with class ID): warplane (n04552348), sports car (n04285008), brambling bird, (n01530575), Siamese cat (n02123597), antelope (n02422699), Swiss mountain dog (n02107574), bull frog (n01641577), garbage truck (n03417042), horse (n02389026), container ship (n03095699).
ImageNet-20 We use the ImageNet-20 in (Ming et al. 2022a), which consists of 20 classes semantically similar to ImageNet-10 (e.g. dog (ID) vs. wolf (OOD)). It contains the following categories: sailboat (n04147183), canoe (n02951358), balloon (n02782093), tank (n04389033), missile (n03773504), bullet train (n02917067), starfish (n02317335), spotted salamander (n01632458), common newt (n01630670), zebra (n01631663), frilled lizard (n02391049), green lizard (n01693334), African crocodile (n01697457), Arctic fox (n02120079), timber wolf (n02114367), brown bear (n02132136), moped (n03785016), steam locomotive (n04310018), space shuttle (n04266014), snowmobile (n04252077).
ImageNet-100 We use the ImageNet-100 in (Ming et al. 2022a), which is curated from ImageNet-1k.
@@ H.2 Datasets
H.3 Sources of model checkpoints
@@ H.4 Hyperparameters
Table 6: Comparison results of CMA with different Temperature parameter $\tau$. The ID dataset is ImageNet-1k.
iNaturalist SUN Places Texture Average
[…]
0.1 21.20 95.94 32.21 93.08 30.33 93.20 49.04 88.11 33.00 92.58
0.2 22.42 95.70 31.10 93.26 29.95 93.20 47.89 88.32 32.49 92.62
0.4 23.22 95.55 30.73 93.34 29.88 93.19 29.88 88.42 28.89 92.63
0.6 23.64 95.49 30.64 93.37 29.88 93.18 47.48 88.45 32.46 92.62
0.8 23.75 95.46 30.55 93.38 29.85 93.18 47.39 88.47 32.42 92.62
1 23.84 95.44 30.51 93.39 29.86 93.17 47.35 88.47 32.41 92.62
1.5 23.99 95.42 30.49 93.40 29.88 93.17 47.32 88.49 32.43 92.62
2 24.02 95.41 30.47 93.40 29.84 93.17 47.26 88.49 32.41 92.62
4 24.14 95.39 30.39 93.41 29.85 93.16 47.25 88.50 32.40 92.62
8 24.21 95.38 30.43 93.41 29.89 93.16 47.23 88.51 32.44 92.62
16 24.23 95.37 30.44 93.41 29.90 93.16 47.23 88.51 32.45 92.61
32 24.29 95.37 30.45 93.42 29.94 93.16 47.26 88.51 32.48 92.62
64 24.27 95.37 30.42 93.42 29.94 93.16 47.25 88.51 32.46 92.62
Table 7: Comparison results of ID classification accuracy on ImageNet-1k (%)
Method ID ACC
zero-shot
MSP (CLIP-B/16) (Hendrycks and Gimpel 2016) 66.78
MSP (CLIP-L/14) (Hendrycks and Gimpel 2016) 73.12
MCM (CLIP-B/16) (Ming et al. 2022a) 67.01
MCM (CLIP-L/14) (Ming et al. 2022a) 73.28
CMA (CLIP-B/16) 66.88
CMA (CLIP-L/14) 73.49
H.5 Prompt-1 to generate Neutral Prompts
Create a series of neutral, generic image prompts that do not specify any particular object, scene, or concept. These prompts should be vague and open-ended, allowing for a wide range of possible interpretations. Follow these guidelines:
1. Start each prompt with "a photo of" or "an image of".
2. Use generic terms like "thing", "object", "item", or "element".
3. Optionally, add a broad context like "in nature", "in a place", or "in an environment".
4. Avoid mentioning any specific objects, living beings, or locations.
5. Keep the prompts simple and concise, typically 5-10 words long.
6. Do not include any adjectives that could narrow down the interpretation.
7. Ensure the prompts are grammatically correct and make sense in English.
8. Create at least 10 unique prompts following these rules.
Examples of the type of prompts you should generate:
- "a photo of a thing in the world"
- "an image of an object somewhere"
- "a picture of something that exists"
- "a photograph of an item in a setting"
Remember, the goal is to create prompts that are as neutral and non-specific as possible while still forming a coherent phrase about an image or photo.
H.6 Prompt-2 to generate Neutral Prompts
Create a series of neutral, abstract image prompts that do not specify any particular object, scene, or concept. These prompts should be vague, open-ended, and focus on general concepts or sensory experiences. Follow these guidelines:
1. Start each prompt with "a photo of", "an image of", or a similar phrase.
2. Use generic terms like "thing", "object", "item", "element", or more abstract concepts like "presence" or "essence".
3. Optionally, add a broad context like "in nature", "in a place", or "in an environment".
4. Avoid mentioning any specific objects, living beings, or locations.
5. Keep the prompts simple and concise, typically 5-15 words long.
6. Do not include any adjectives that could narrow down the interpretation.
7. Ensure the prompts are grammatically correct and make sense in English.
8. Focus on general concepts rather than specific items.
9. Use neutral descriptive language that evokes feelings or imagery without defining what is being referred to.
10. Include references to sensory experiences (sight, sound, touch, smell, taste) without attributing them to particular phenomena.
11. Use broad terms instead of elements that could be interpreted as belonging to a particular category.
12. Employ indefinite articles (a/an) and vague quantifiers (some, many) to maintain ambiguity.
13. Create at least 10 unique prompts following these rules.
Examples of the type of prompts you should generate:
- "a photo of a thing in the world"
- "an image of an abstract presence in a space"
- "a picture of something evoking a gentle sensation"
- "a photograph of an undefined texture in an environment"
- "an image capturing a vague impression of color"
Remember, the goal is to create prompts that are as neutral, abstract, and non-specific as possible while still forming a coherent phrase about an image or photo. Focus on sensory elements and general concepts without implying any specific meaning or object.